a plane can fly 200 miles against the wind in the same amount of time as it can fly 300 miles with the wind. if the velocity of the wind is 40 mph, what is the rate, r, of plane in still air?
Answer by richwmiller(17219)

r*t=d
t=d/r
200/(v-40)=300/(v+40)
200*(v+40)=300(v-40)
200v+8000=300v-12000
20000=100v
200=v